Jefferson Lab HALL A Logbook for April 07 - All EXCEPT Automatic Entries

  • Main INDEX

    Monday

  • 200669 04/30/07 22:46 K.Allada Target moved to dummy
  • 200655 04/30/07 19:32 bogdanw radiation monitor
  • 200652 04/30/07 19:09 Jack Raster work during earlier access
  • 200642 04/30/07 16:50 bogdanw HV
  • 200641 04/30/07 16:45 bogdanw Hall A Tools
  • 200640 04/30/07 16:44 K.Allada raster size - spot++
  • 200636 04/30/07 16:00 X. Zheng day shift summary
  • 200635 04/30/07 14:24 jpchen/gilma cryotarget (He4) cool-down completed
  • 200634 04/30/07 14:05 X. Zheng Hall update
  • 200633 04/30/07 12:40 gilman cryotarget vented
  • 200632 04/30/07 12:32 X. Zheng target warm up
  • 200631 04/30/07 12:06 M. Stevens ESR trip
  • 200630 04/30/07 11:48 X. Zheng more alarms
  • 200629 04/30/07 11:42 X. Zheng lost right Q3
  • 200627 04/30/07 11:39 gilman esr 4k flow unstable
  • 200608 04/30/07 08:01 Moffit Owl Shift Summary
  • 200601 04/30/07 06:38 Elena Kh. He4 target
  • 200599 04/30/07 05:58 Moffit Shift screen captures.. a mixed assortment.
  • 200596 04/30/07 05:25 Moffit Raster is rastering on Al dummy.
  • 200594 04/30/07 05:06 Elena Kh. Target screens
  • 200593 04/30/07 05:05 Elena Kh. Dummy target
  • 200585 04/30/07 03:23 Moffit Raster is still rastering.
  • 200574 04/30/07 00:56 Moffit Late Entry: Target OTR removed
  • Sunday

  • 200568 04/29/07 23:55 K.Allada Shift/Run summary
  • 200557 04/29/07 21:43 K.Allada target OTR
  • 200556 04/29/07 21:41 kelleher target screens
  • 200554 04/29/07 21:05 K.Allada back to production with 4He
  • 200552 04/29/07 21:02 K.Allada Raster during dummy run
  • 200549 04/29/07 20:50 K.Allada Hall A Tools screen shot
  • 200542 04/29/07 17:40 K.Allada Radiation monitor strip chart
  • 200537 04/29/07 16:39 K.Allada Raster size - spot++
  • 200531 04/29/07 16:04 N. Sparveris Shift Summary
  • 200530 04/29/07 15:47 gomez General Accelerator Parameters - 4/29/07 -15:47
  • 200529 04/29/07 15:36 gomez Target Status 2 - 4/29/07 - 15:36
  • 200526 04/29/07 15:21 gomez Target status 1 - 4/29/07 - 15:20
  • 200525 04/29/07 15:18 gomez Right detector HV - 4/29/07 - 15:17
  • 200524 04/29/07 15:17 gomez Left detector HV - 4/29/07 - 15:16
  • 200523 04/29/07 15:15 gomez Hall A Tools & BCM - 4/29/07 - 15:15
  • 200522 04/29/07 15:14 gomez Rad. Monitors & BCM - 4/49/07 - 15:12
  • 200517 04/29/07 14:09 gomez run#2982 - spot++
  • 200512 04/29/07 13:23 gomez run#2980 -spot++
  • 200503 04/29/07 11:25 N. Sparveris Radiation Monitors strip-chart
  • 200500 04/29/07 10:44 Elena Kh. Target screens
  • 200499 04/29/07 10:41 N. Sparveris Hall A Tools Screen Capture
  • 200486 04/29/07 08:04 V. Sulkosky Owl Shift Summary
  • 200482 04/29/07 07:33 V. Sulkosky Late Target Entry
  • 200481 04/29/07 07:29 V. Sulkosky BCM and Radiation Monitor Charts
  • 200474 04/29/07 06:02 V. Sulkosky Hall A Tools Snapshot
  • 200473 04/29/07 05:52 V. Sulkosky Raster Check, Run 2965
  • 200468 04/29/07 05:12 R. Michaels re:raster repair
  • 200462 04/29/07 04:44 R. Michaels raster repair (but for how long ?)
  • 200461 04/29/07 02:51 R. Michaels plan to replace raster driver, I'm coming in .....
  • 200460 04/29/07 02:21 V. Sulkosky Hall A Tools Screen
  • 200459 04/29/07 02:20 V. Sulkosky Raster Woes.
  • 200456 04/29/07 01:45 V. Sulkosky OTR Profile on BeO Target
  • 200454 04/29/07 01:16 Mina Kat spot++ run 2959 old raster size
  • 200451 04/29/07 01:05 Mina Kat spot++ run 2958
  • 200450 04/29/07 01:03 V. Sulkosky OTR Profile
  • 200446 04/29/07 00:34 V. Sulkosky Raster Repaired
  • 200445 04/29/07 00:09 Makis P Shift Summary
  • Saturday

  • 200443 04/28/07 23:56 Makis P Hall Access
  • 200442 04/28/07 23:38 Makis P Hall A access request
  • 200441 04/28/07 23:32 Makis P Reducing raster size by 20%
  • 200440 04/28/07 23:28 Makis P Vertical RASTER died again
  • 200436 04/28/07 22:34 Mina Kat Target controls screen
  • 200431 04/28/07 21:44 Mina Kat spot++ run 2953
  • 200430 04/28/07 21:43 Makis P Radiation Monitors strip-chart
  • 200429 04/28/07 21:42 Makis P Current Monitor Run 2953
  • 200428 04/28/07 21:41 Makis p Hall A Tools Run 2953
  • 200419 04/28/07 20:07 Makis P Beam Way VESTA alert
  • 200414 04/28/07 19:00 Mina Kat Target charts
  • 200405 04/28/07 16:57 Mina Kat spot+= run 2943
  • 200400 04/28/07 16:05 X. Zheng day shift summary
  • 200393 04/28/07 14:52 X. Zheng radiation monitor stripchart
  • 200387 04/28/07 13:32 X. Zheng Hall A Tools Screen Capture
  • 200386 04/28/07 13:32 X. Zheng spot++
  • 200384 04/28/07 13:22 X. Zheng confusion about the Moller target position
  • 200383 04/28/07 12:41 X. Zheng raster fixed
  • 200382 04/28/07 10:59 X. Zheng controlled access
  • 200381 04/28/07 10:27 X. Zheng raster update
  • 200380 04/28/07 09:50 X. Zheng lost raster magnet
  • 200378 04/28/07 09:37 X. Zheng spot++ (good)
  • 200377 04/28/07 09:35 X. Zheng choppy beam
  • 200369 04/28/07 08:04 V. Sulkosky Owl Shift Summary
  • 200365 04/28/07 07:34 Elena Kh. He4 target
  • 200364 04/28/07 07:26 V. Sulkosky Radiation Monitor Snapshot
  • 200362 04/28/07 07:25 Elena Kh. runs:2926 - 2930 are with dummy target
  • 200357 04/28/07 05:51 V. Sulkosky Hall A Tools Snapshot
  • 200354 04/28/07 05:33 V. Sulkosky BPMA Y Position Off during Run 2926
  • 200353 04/28/07 05:28 Elena Kh. spot++
  • 200349 04/28/07 04:41 Elena Kh. Dummy 20cm target
  • 200347 04/28/07 04:02 Elena Kh. Target screens
  • 200342 04/28/07 03:22 V. Sulkosky Radiation Monitor and BCM Strip Charts
  • 200339 04/28/07 02:52 V. Sulkosky No Beam
  • 200336 04/28/07 02:07 V. Sulkosky Hall A Tools Screen
  • 200331 04/28/07 01:23 V. Sulkosky Raster Captured
  • 200325 04/28/07 00:05 johna Swing Shift Summary
  • Friday

  • 200317 04/27/07 21:55 johna MCC calls; no beam for 15 min.
  • 200313 04/27/07 21:46 johna BCM and radiation monitor strip charts
  • 200306 04/27/07 20:40 johna Hall A Tools Screen
  • 200303 04/27/07 20:04 Mina Kat Target Charts
  • 200300 04/27/07 19:45 johna Beam is back
  • 200298 04/27/07 19:17 Mina Kat Moved target to He4
  • 200297 04/27/07 18:44 Mina Kat Moved to empty target
  • 200296 04/27/07 18:17 johna Injector problems
  • 200295 04/27/07 16:31 johna Recovered from magnet problems, trying to get beam. MCC having injector problems
  • 200294 04/27/07 16:01 X. Zheng day shift summary
  • 200292 04/27/07 15:26 dezern lead flow-magnet quench
  • 200291 04/27/07 14:12 X. Zheng controlled access
  • 200290 04/27/07 13:49 X. Zheng TECH - cryo alarm - lost both spectrometer
  • 200289 04/27/07 12:27 X. Zheng radiation monitor stripchart
  • 200288 04/27/07 12:27 X. Zheng Beam current monitor
  • 200278 04/27/07 10:42 X. Zheng spot++
  • 200276 04/27/07 09:02 X. Zheng MCC update
  • 200275 04/27/07 08:59 X. Zheng RF problem, no beam
  • 200269 04/27/07 08:10 kelleher Owl Shift Summary
  • 200262 04/27/07 06:57 kelleher radiation monitors
  • 200260 04/27/07 06:25 kelleher hall a tool screen
  • 200250 04/27/07 05:06 kelleher raster scan after down
  • 200247 04/27/07 04:37 kelleher raster fixed
  • 200246 04/27/07 03:39 kelleher raster magnet down
  • 200242 04/27/07 02:59 kelleher raster capture
  • 200236 04/27/07 02:14 atrig Trigger-Download
  • 200232 04/27/07 01:37 kelleher right S2 scaler fixed
  • 200227 04/27/07 01:04 Chiranjib TARGET SCREENS II
  • 200226 04/27/07 01:02 Chiranjib TAGET SCREENS I
  • 200225 04/27/07 01:00 Chiranjib CODA rebooted
  • 200221 04/27/07 00:50 kelleher right arm S2M rates
  • Thursday

  • 200201 04/26/07 21:56 Margaziotis Target moved
  • 200198 04/26/07 21:03 johna Beam down for ~10 min
  • 200196 04/26/07 20:48 johna spot++
  • 200195 04/26/07 20:43 johna radiation strip chart
  • 200193 04/26/07 20:42 johna Taking dummy target data
  • 200189 04/26/07 19:34 johna Left arm Q3 drifting a little
  • 200184 04/26/07 18:32 johna Hall A General Tools Screen and BCM stripchart
  • 200181 04/26/07 18:03 folts target collimator
  • 200174 04/26/07 17:05 johna Q3 drift at end of run 2859
  • 200164 04/26/07 15:59 Higinbotham Shift Summary
  • 200160 04/26/07 15:24 Higinbotham Problem Fixed
  • 200159 04/26/07 15:08 Higinbotham Lost Right Arm IOC @ 3pm
  • 200149 04/26/07 13:30 Elena Kh. vdc hv -> start of run
  • 200142 04/26/07 13:06 Higinbotham Increase VDC Voltage
  • 200135 04/26/07 12:08 Higinbotham Spot Check
  • 200133 04/26/07 11:35 Higinbotham Tools Screen
  • 200132 04/26/07 11:35 Higinbotham Changed to Q2=55m-2
  • 200131 04/26/07 09:17 atrig Trigger-Download
  • 200128 04/26/07 07:53 kelleher radiation monitor stripchart
  • 200124 04/26/07 07:21 xqian target status
  • 200123 04/26/07 07:20 xqian target movement
  • 200121 04/26/07 07:05 kelleher plot of raster
  • 200118 04/26/07 06:53 kelleher Hall A Tools
  • 200107 04/26/07 05:13 xqian target alarm
  • 200092 04/26/07 03:37 kelleher 4He Cell Collimator control picture
  • Wednesday

  • 200056 04/25/07 22:28 johna radiation monitor strip chart
  • 200049 04/25/07 21:43 Margaziotis Cryostat/ESR screens
  • 200048 04/25/07 21:39 Margaziotis Target screens
  • 200045 04/25/07 21:21 johna BCM stripchart
  • 200044 04/25/07 21:16 johna General Tools Screen
  • 200043 04/25/07 21:15 johna spot++ screen
  • 200031 04/25/07 19:21 Margaziotis Target changes
  • 200027 04/25/07 17:54 R. Michaels stability of timing offset T0
  • 200025 04/25/07 17:23 folts Reminder to all shift workers
  • 200023 04/25/07 17:16 R. Michaels spot++
  • 200019 04/25/07 16:24 johna Unsual spot++ result
  • 200017 04/25/07 15:55 L. Zhu Day Shift Summary
  • 200006 04/25/07 14:15 Dezern/Fansl Magnet Lead Flows
  • 199997 04/25/07 13:02 B Norum Magnet
  • 199995 04/25/07 12:57 L. Zhu Hall A tools
  • 199991 04/25/07 10:44 R. Michaels swapped CAMAC delay to fix missing TDC
  • 199989 04/25/07 10:31 atrig Trigger-Download
  • 199988 04/25/07 10:23 L. Zhu controlled access
  • 199987 04/25/07 10:08 Jack Left Shield House AHU
  • 199986 04/25/07 09:42 R. Michaels Missing TDC
  • 199985 04/25/07 09:28 L. Zhu magnets gone
  • 199984 04/25/07 09:15 B. Norum LHRS, magnet trip
  • 199978 04/25/07 08:01 A. Puckett requested controlled access for AC in LHRS
  • 199974 04/25/07 07:31 A. Puckett back on 4He, spot check
  • 199973 04/25/07 07:30 A. Puckett Back to 4He
  • 199971 04/25/07 07:19 xqian target movement
  • 199969 04/25/07 06:45 xqian target status
  • 199968 04/25/07 06:45 xqian target movement
  • 199967 04/25/07 06:19 A. Puckett tools and BPM screengrab
  • 199966 04/25/07 06:17 A. Puckett a spot check
  • 199960 04/25/07 04:11 A. Puckett changed to dummy target
  • 199948 04/25/07 02:31 A. Puckett But, left S1 TDC channel still missing
  • 199947 04/25/07 02:20 A. Puckett Left shield hut temp. looks stable.
  • 199943 04/25/07 02:03 A. Camsonne Re : Temperature in shielding hut
  • 199941 04/25/07 01:52 A. Puckett Left Arm shield hut temp.
  • 199938 04/25/07 01:39 A. Puckett 4He Production statistics so far...
  • 199929 04/25/07 00:16 johna Shift summary
  • Tuesday

  • 199921 04/24/07 23:18 A. Camsonne Delay module channel going bad ?
  • 199916 04/24/07 22:34 johna TDC channel is missing again
  • 199915 04/24/07 22:32 johna Hall A Tools screen
  • 199914 04/24/07 22:29 johna spot++ for run 2760
  • 199910 04/24/07 21:01 johna Beam down - problem in injector
  • 199905 04/24/07 19:57 Jack Left dipole fault
  • 199903 04/24/07 18:50 R. Feuerbach bad cable in L-S1 TDC signal path
  • 199902 04/24/07 18:46 johna tools window showing He flow problems
  • 199901 04/24/07 18:38 johna Missing TDC is back, but He flow to magnets went away
  • 199898 04/24/07 18:22 johna Missing TDC: Left arm - S1plane, 5th paddle, right side
  • 199897 04/24/07 18:18 atrig Trigger-Download
  • 199883 04/24/07 16:00 kuchina Day shift summary
  • 199878 04/24/07 15:11 Jack Ozone monitor display now in ppm
  • 199876 04/24/07 14:44 Ole Hansen Late entry: spot++ for run 2739
  • 199874 04/24/07 14:30 Norum Beam
  • 199867 04/24/07 13:02 kuchina Current Monitor screen
  • 199866 04/24/07 13:00 kuchina General Tools screen
  • 199863 04/24/07 12:12 Ole Hansen No obvious evidence of dummy target damage
  • 199862 04/24/07 12:08 Ole Hansen spot++ run 2740 on 4He
  • 199859 04/24/07 11:22 Ole Hansen Run 2739 - test run with bigger raster
  • 199857 04/24/07 11:21 Ole Hansen Overcurrent on dummy target
  • 199855 04/24/07 11:13 kuchina Q1 is back
  • 199854 04/24/07 11:12 Ole Hansen spot++ run 2736
  • 199853 04/24/07 11:10 Ole Hansen spot++ run 2735
  • 199852 04/24/07 11:09 Ole Hansen spot++ run 2737
  • 199850 04/24/07 10:50 kuchina Left Q1 tripped
  • 199848 04/24/07 10:47 Ole Hansen Run 2737 - magnet crash
  • 199846 04/24/07 10:39 kuchina Right Magnet Alarm
  • 199844 04/24/07 09:59 R. Michaels trig. stability check (good)
  • 199840 04/24/07 09:06 Ole Hansen Target to dummy
  • 199831 04/24/07 08:02 A. Puckett Owl shift summary
  • 199830 04/24/07 07:56 xqian Target movement
  • 199829 04/24/07 07:55 xqian Target Status
  • 199800 04/24/07 03:30 A. Puckett running on 4He now
  • 199796 04/24/07 02:59 A. Puckett spot++ at 80 muA on BeO
  • 199790 04/24/07 02:11 A. Puckett Raster Problems, Solved
  • 199785 04/24/07 01:28 A. Puckett Raster Problems, cont.
  • 199784 04/24/07 01:13 A. Puckett Raster problems, cont.
  • 199782 04/24/07 01:01 A. Puckett Raster Problem...
  • Monday

  • 199779 04/23/07 22:29 johna MCC calls - trying to deliver beam, but moller target was in
  • 199778 04/23/07 22:01 Markowitz Empty target in
  • 199777 04/23/07 18:08 johna Sweep in progress
  • 199776 04/23/07 16:32 Elena Kh. 50 Q2 kinematics
  • 199775 04/23/07 16:21 jpchen He4 target cooldown completed
  • 199774 04/23/07 16:16 atrig Trigger-Download
  • 199773 04/23/07 16:13 Elena Kh. R VDC HV -> 3.6kV
  • 199772 04/23/07 16:11 Elena Kh. R S1 HVs -> 75%
  • 199771 04/23/07 15:18 spiegel Pre-beam checklist
  • 199761 04/23/07 12:23 R. Michaels T5, T6 overlap check
  • 199760 04/23/07 12:12 atrig Trigger-Download
  • 199759 04/23/07 12:11 atrig Trigger-Download
  • 199758 04/23/07 12:10 atrig Trigger-Download
  • 199757 04/23/07 12:09 atrig Trigger-Download
  • 199756 04/23/07 11:46 atrig Trigger-Download
  • 199755 04/23/07 11:45 atrig Trigger-Download
  • 199754 04/23/07 11:45 Fansler/Stev R Q1 Lead Flow Offset
  • 199753 04/23/07 11:44 atrig Trigger-Download
  • 199752 04/23/07 11:44 atrig Trigger-Download
  • 199751 04/23/07 11:39 atrig Trigger-Download
  • 199750 04/23/07 11:39 atrig Trigger-Download
  • 199749 04/23/07 11:34 atrig Trigger-Download
  • 199748 04/23/07 11:33 atrig Trigger-Download
  • 199747 04/23/07 11:29 atrig Trigger-Download
  • 199746 04/23/07 11:29 atrig Trigger-Download
  • 199745 04/23/07 11:28 atrig Trigger-Download
  • 199744 04/23/07 11:27 atrig Trigger-Download
  • 199743 04/23/07 11:26 atrig Trigger-Download
  • 199742 04/23/07 11:26 atrig Trigger-Download
  • 199741 04/23/07 11:07 R. Michaels timing checks for T5 and T6
  • 199740 04/23/07 10:52 atrig Trigger-Download
  • 199739 04/23/07 10:52 atrig Trigger-Download
  • 199738 04/23/07 10:51 atrig Trigger-Download
  • 199733 04/23/07 10:45 atrig Trigger-Download
  • 199732 04/23/07 10:44 atrig Trigger-Download
  • 199731 04/23/07 10:43 atrig Trigger-Download
  • 199730 04/23/07 10:38 atrig Trigger-Download
  • 199728 04/23/07 10:37 atrig Trigger-Download
  • 199723 04/23/07 10:34 atrig Trigger-Download
  • 199721 04/23/07 10:31 atrig Trigger-Download
  • 199720 04/23/07 10:30 atrig Trigger-Download
  • 199718 04/23/07 10:22 atrig Trigger-Download
  • Friday

  • 199717 04/20/07 16:33 fansler R Q1 Lead flow A
  • 199716 04/20/07 16:32 dezern beam line girder turbo
  • 199713 04/20/07 13:31 folts smoke removal fans
  • 199712 04/20/07 13:21 folts Hall status
  • 199708 04/20/07 09:59 jpchen Ozone stripe chart
  • 199707 04/20/07 09:47 jpchen target warm up completed
  • 199706 04/20/07 09:16 jpchen Hall A in restricted access, please contact Ed/JP before starting any work in the hall
  • 199705 04/20/07 09:10 Albert/JP beam stopped. power down magents, target warm
  • 199704 04/20/07 09:03 Shahinyan Target status
  • 199702 04/20/07 07:52 kelleher slowly closing JT valve
  • 199701 04/20/07 07:51 R. Michaels owl shift summary
  • 199697 04/20/07 06:59 R. Michaels (minor) halla_scaler_process error
  • 199694 04/20/07 06:42 kelleher target screens
  • 199688 04/20/07 05:30 kelleher target motion: 4He
  • 199685 04/20/07 04:10 kelleher target motion: Al dummy cell
  • 199684 04/20/07 04:04 R. Michaels move target to Al dummy
  • 199683 04/20/07 04:01 R. Michaels spot++
  • 199673 04/20/07 01:53 R. Michaels Solved EPICS problems.
  • 199670 04/20/07 01:42 R. Michaels reboot iocha5 (BCM)
  • 199665 04/20/07 00:38 R. Michaels gen. tools
  • 199662 04/20/07 00:02 Solvignon Shift summary
  • 199660 04/20/07 00:01 Solvignon Ozone stripchart
  • Thursday

  • 199657 04/19/07 23:51 Solvignon Hall A general tools for Q2=50fm-2
  • 199655 04/19/07 23:41 Solvignon Right VDC HV changed
  • 199654 04/19/07 23:10 atrig Trigger-Download
  • 199649 04/19/07 22:24 Markowitz move to dummy target
  • 199645 04/19/07 21:45 Solvignon New beam current limit: 85uA
  • 199644 04/19/07 21:32 Markowitz target screens
  • 199641 04/19/07 21:18 Solvignon Hall A general tools
  • 199640 04/19/07 21:13 Solvignon Spot++
  • 199637 04/19/07 21:01 Solvignon LQ3 went slightly off then LDipole went to wrong value
  • 199631 04/19/07 19:54 Solvignon Right arm VDC HV raised to 3.6kV
  • 199630 04/19/07 19:53 Markowitz opened JT valve by .25 step
  • 199628 04/19/07 19:34 Jack Ozone monitor stripchart
  • 199625 04/19/07 19:26 Mina Kat Changed hist limits for raster vs current in spot++
  • 199620 04/19/07 18:26 folts NEW HALL ENTRY REQUIRMENTS
  • 199610 04/19/07 16:49 jpchen lowered target T to 8.5K.
  • 199606 04/19/07 16:09 Higinbotham Kinematics 45 fm-2
  • 199605 04/19/07 16:02 Markowitz Trigger downloaded
  • 199604 04/19/07 16:00 Higinbotham Shift Summary
  • 199603 04/19/07 16:00 atrig Trigger-Download
  • 199602 04/19/07 15:29 Higinbotham Right S1 to 75%
  • 199601 04/19/07 15:22 L. Zhu target screens
  • 199584 04/19/07 11:46 Higinbotham ARC PMT Ch.1 Problem!
  • 199583 04/19/07 11:22 jennifer air sampling 4/19/07
  • 199582 04/19/07 11:17 Higinbotham S1 changed to 70% from 60%
  • 199581 04/19/07 10:44 jpchen target T to 8.7K
  • 199580 04/19/07 10:13 Jack Late entry/techs Right Q1
  • 199576 04/19/07 09:28 L. Zhu lower events limit
  • 199568 04/19/07 07:58 A. Puckett Owl shift summary
  • 199567 04/19/07 07:45 kelleher target screens
  • 199565 04/19/07 07:27 kelleher target moved to 4He
  • 199563 04/19/07 07:00 A. Puckett moved LHRS collimator to 6 msr
  • 199554 04/19/07 04:56 A. Puckett rad. monitor stripcharts
  • 199551 04/19/07 04:45 A. Puckett BPM stripchart
  • 199550 04/19/07 04:45 A. Puckett tools and BCM screengrab
  • 199541 04/19/07 03:38 A. Puckett "final" spot size
  • 199537 04/19/07 03:24 kelleher target to 4He
  • 199536 04/19/07 03:20 A. Puckett spot++ on BeO after beam restore
  • 199518 04/19/07 02:15 A. Puckett MCC can't get HARPs working
  • Wednesday

  • 199514 04/18/07 23:56 Solvignon Shift summary
  • 199513 04/18/07 23:54 atrig Trigger-Download
  • 199512 04/18/07 18:08 folts late entry/ tech
  • 199509 04/18/07 18:04 atrig Trigger-Download
  • 199508 04/18/07 18:02 atrig Trigger-Download
  • 199507 04/18/07 17:53 Elena Kh. Right s1 HVs -> 64%
  • 199506 04/18/07 17:48 Elena Kh. MBSY1C Bdl ?
  • 199504 04/18/07 15:50 L. Zhu day shift summary
  • 199503 04/18/07 15:50 Elena Kh. Updated db_L.prl1.dat
  • 199502 04/18/07 15:49 L. Zhu Hall A Tools
  • 199501 04/18/07 14:23 jpchen 4He operation condition.
  • 199500 04/18/07 13:44 jennifer air sampling
  • 199499 04/18/07 12:54 L. Zhu target screens for he4
  • 199498 04/18/07 12:36 sawatzky File split problem
  • 199497 04/18/07 12:28 R. Michaels target change complete (He4)
  • 199496 04/18/07 10:19 R. Michaels tools
  • 199494 04/18/07 10:03 R. Michaels trying to use 4 GB filesplit (failed)
  • 199493 04/18/07 09:56 R. Michaels pedrun, restored peds
  • 199488 04/18/07 09:34 R. Michaels down for energy/tgt change
  • 199483 04/18/07 08:55 L. Zhu start kinematics change
  • 199478 04/18/07 08:10 R. Michaels target GUIs
  • 199474 04/18/07 08:00 A. Puckett Owl Shift Summary
  • 199467 04/18/07 06:18 A. Puckett spot++
  • 199466 04/18/07 06:17 A. Puckett rad. mon stripchart
  • 199465 04/18/07 06:16 A. Puckett BPM stripchart
  • 199463 04/18/07 06:15 A. Puckett BCM and Hall A Tools Screen Grab
  • 199461 04/18/07 05:47 A. Puckett CODA crashed at end of run 2613, TS11 rebooted
  • 199453 04/18/07 04:16 Elena Kh. Target screens
  • 199452 04/18/07 03:58 A. Puckett CODA problems
  • 199445 04/18/07 02:03 Elena Kh. spot++
  • 199435 04/18/07 00:00 Solvignon Shift summary
  • Tuesday

  • 199426 04/17/07 22:13 Chiranjib TARGET SCREENS II
  • 199425 04/17/07 22:12 Chiranjib TARGET SCREENS I
  • 199419 04/17/07 21:19 Solvignon Radiation monitors stripcharts
  • 199418 04/17/07 21:18 Solvignon Screenshot of Hall A general tools
  • 199416 04/17/07 21:12 Solvignon Spot++
  • 199409 04/17/07 19:46 Chiranjib TARGET SCREENS II
  • 199408 04/17/07 19:45 Chiranjib TARGET SCREENS I
  • 199405 04/17/07 19:15 Solvignon 1 hour at 100uA without beamtrip
  • 199398 04/17/07 18:03 Chiranjib CODA working now
  • 199395 04/17/07 17:34 Chiranjib Components for RUN CONTROL rebooted
  • 199392 04/17/07 17:27 Chiranjib RUN control failure
  • 199388 04/17/07 17:07 Solvignon Running at 100uA
  • 199387 04/17/07 17:01 TECHS RIGHT Q1 COMM ERROR/EARTH LEAKAGE FAILURE
  • 199385 04/17/07 16:39 jpchen right Q1 trip fixed by Hall A tech's
  • 199384 04/17/07 16:02 kuchina shift summary
  • 199382 04/17/07 15:30 kuchina Right Q1 trip
  • 199378 04/17/07 15:12 kuchina Hall A Current Monitor
  • 199377 04/17/07 15:11 kuchina General tools screenshot
  • 199358 04/17/07 09:49 Shahinyan Stop run
  • 199356 04/17/07 09:20 Shahinyan Target status
  • 199355 04/17/07 09:07 Kuchina spot++
  • 199351 04/17/07 08:01 A. Puckett Owl Shift Summary
  • 199335 04/17/07 04:50 A. Puckett rad monitors after beam restore
  • 199334 04/17/07 04:45 A. Puckett spot check after beam restored
  • 199332 04/17/07 04:33 A. Puckett RF Cavity problem, beam down
  • 199330 04/17/07 04:10 Elena Kh. spot++
  • 199327 04/17/07 03:58 Elena Kh. spot++
  • 199324 04/17/07 03:23 Elena Kh. Target screens
  • 199319 04/17/07 02:20 Elena Kh. spot++
  • 199316 04/17/07 01:57 Elena Kh. Changed pedestal suppression threshold before run 2559
  • 199315 04/17/07 01:53 A. Puckett Rad. mon stripcharts:
  • 199313 04/17/07 01:46 A. Puckett MCC performs QE measurement.
  • 199312 04/17/07 01:41 A. Puckett spot check during run 2556
  • 199303 04/17/07 00:01 Solvignon Shift summary
  • Monday

  • 199298 04/16/07 23:18 Chiranjib TARGET SCREENS 2
  • 199297 04/16/07 23:17 Chiranjib TARGET SCREENS 1
  • 199288 04/16/07 21:19 solvignon radiation monitor stripcharts
  • 199287 04/16/07 21:03 Chiranjib Target screens 2
  • 199286 04/16/07 21:03 Chiranjib target screens
  • 199279 04/16/07 19:47 Solvignon Spot++ for 1st production run of this evening
  • 199278 04/16/07 19:44 Solvignon Hall A general tools
  • 199276 04/16/07 19:26 Chiranjib Target --> He-3
  • 199274 04/16/07 19:19 Solvignon Beam spot on BeO and camera
  • 199272 04/16/07 18:53 Solvignon Beam back
  • 199271 04/16/07 18:51 Solvignon Controlled access done
  • 199270 04/16/07 18:43 M. Stevens Left Q1 spurious trips
  • 199269 04/16/07 16:19 sawatzky Redownload of trigger just to be safe
  • 199268 04/16/07 16:17 sawatzky Shift Summary DAY -- Monday Apr. 16
  • 199267 04/16/07 16:14 atrig Trigger-Download
  • 199261 04/16/07 13:36 sawatzky Hall A FFB system has been modified by MCC
  • 199260 04/16/07 13:30 sawatzky Radiation strip charts
  • 199259 04/16/07 11:17 sawatzky Gas panel alarm followup
  • 199258 04/16/07 11:10 sawatzky Gas System Interlock alarm (intermittent)
  • 199257 04/16/07 10:55 sawatzky Going to Controlled Access
  • 199256 04/16/07 10:55 sawatzky Analyzer hypothesis update
  • 199253 04/16/07 10:09 sawatzky Analyzer replay problem hypothesis
  • 199252 04/16/07 09:51 sawatzky Left Q1 tripped
  • 199251 04/16/07 09:11 sawatzky Beam studies begins
  • 199250 04/16/07 09:06 L. Zhu more target screens
  • 199246 04/16/07 08:58 L. Zhu target screens
  • 199243 04/16/07 08:06 Elena Kh. Shift Summary
  • 199239 04/16/07 07:49 Elena Kh. spot++
  • 199236 04/16/07 07:15 Elena Kh. no left vdc eff in run 2535 again
  • 199235 04/16/07 07:07 Elena Kh. replay ?s
  • 199230 04/16/07 06:08 Elena Kh. Current -> 95uA
  • 199219 04/16/07 03:34 Elena Kh. Changed ped sub -> 10 for Ls1,s2 and Rs1,s2
  • 199216 04/16/07 02:59 Elena Kh. Radiation monitors
  • 199215 04/16/07 02:58 Elena Kh. Hall A Genelar tools
  • 199214 04/16/07 02:56 Elena Kh. beam_info.stp and current monitor
  • 199213 04/16/07 02:45 Elena Kh. Current -> 100uA
  • 199210 04/16/07 02:09 Elena Kh. ped cuts on shower -> 10 from run 2527
  • 199207 04/16/07 01:33 gomez Gas Shed Airflow alarm
  • 199206 04/16/07 01:18 Elena Kh. Run 2525 without end: ROC2 fail
  • Sunday

  • 199199 04/15/07 23:59 kelleher Swing Shift Summary
  • 199196 04/15/07 23:41 kelleher raster scan for run #2522
  • 199195 04/15/07 23:25 Kalyan CODA crashed
  • 199191 04/15/07 21:58 kelleher radiation strip chart
  • 199188 04/15/07 21:34 Kalyan Target screens
  • 199186 04/15/07 21:23 Kelleher Left Q1 trip
  • 199182 04/15/07 20:55 Kalyan back to 95uA beam current
  • 199179 04/15/07 20:14 Kalyan Beam current reduced to 93uA
  • 199176 04/15/07 20:03 kelleher raster scan for run 2515
  • 199164 04/15/07 16:00 Solvignon Shift summary
  • 199162 04/15/07 15:45 xin Target Status
  • 199161 04/15/07 15:16 Solvignon Radiation monitor stripcharts
  • 199160 04/15/07 15:15 Solvignon Hall A general tools screenshot
  • 199156 04/15/07 13:58 Solvignon Beam spot
  • 199150 04/15/07 13:31 Solvignon Beam back
  • 199149 04/15/07 13:02 xin reboot ROC3
  • 199148 04/15/07 13:00 Solvignon No beam for the next 30 to 45 minutes
  • 199143 04/15/07 12:04 Solvignon Left Q1 tripped
  • 199137 04/15/07 10:15 Solvignon Beam halo at the beginning of run 2498
  • 199127 04/15/07 08:03 Elena Kh. Shift Summary
  • 199126 04/15/07 07:54 Elena Kh. R VDC HVs -> 3.5kV
  • 199119 04/15/07 06:02 Elena Kh. spot++
  • 199116 04/15/07 05:23 Elena Kh. spot++
  • 199115 04/15/07 05:12 Elena Kh. run 2490
  • 199112 04/15/07 05:09 Elena Kh. Current -> 80uA
  • 199105 04/15/07 03:46 Elena Kh. Radiation monitors stripcharts
  • 199104 04/15/07 03:45 Elena Kh. Hall A General Tools and Current Monitor
  • 199103 04/15/07 03:43 Elena Kh. beam info stripchart
  • 199102 04/15/07 03:41 Elena Kh. spot++
  • 199099 04/15/07 03:10 Elena Kh. changed db_run.dat update
  • 199098 04/15/07 02:40 Elena Kh. Left Q1 trip
  • 199095 04/15/07 02:28 Elena Kh. R VDC hits and eff for run 2485
  • 199092 04/15/07 01:35 Elena Kh. Changed R VDC HVs ->3.55 kV
  • 199089 04/15/07 01:21 Elena Kh. spot++
  • 199088 04/15/07 01:18 Elena Kh. Current -> 85uA
  • 199080 04/15/07 00:04 Mina Kat Shift Summary
  • Saturday

  • 199079 04/14/07 23:55 Makis P Full analyses checks
  • 199078 04/14/07 23:35 Mina Kat Left Q1 trip
  • 199075 04/14/07 23:08 Makis P Left v2 VDC: one noisy and one dead wire
  • 199072 04/14/07 23:02 A. Camsonne End of run left arm scalers
  • 199068 04/14/07 21:50 Mina Kat Frequent beam trips
  • 199066 04/14/07 21:45 Mina Kat spot++ run 2476
  • 199065 04/14/07 21:40 Makis P Target controls screens
  • 199064 04/14/07 21:38 Makis P Target T, P etc charts
  • 199059 04/14/07 20:40 Mina Kat spot++ run 2474
  • 199058 04/14/07 20:26 Mina Kat Left Q1 trip
  • 199055 04/14/07 20:14 Mina Kat general Tools and current monitor
  • 199052 04/14/07 19:54 Mina Kat rad monitors strip tools
  • 199051 04/14/07 19:48 Mina Kat VESDA Beam Way alert
  • 199046 04/14/07 19:05 Mina Kat Expected current for later on tonight
  • 199043 04/14/07 18:17 Mina Kat Beam ifo strip tool
  • 199038 04/14/07 16:53 Mina Kat spot++ run 2647
  • 199035 04/14/07 16:15 Elena,Javier Shift Summary
  • 199031 04/14/07 15:46 xin Target status
  • 199028 04/14/07 15:19 gomez spot - run# 2464
  • 199023 04/14/07 14:15 gomez spot++ 2462
  • 199022 04/14/07 14:11 gomez Left Q1 crashed - run 2461
  • 199019 04/14/07 14:00 gomez VESDA alert - 2nd time
  • 199018 04/14/07 13:45 gomez spot++ for run# 2461
  • 199017 04/14/07 13:40 Jack VESDA alerts
  • 199013 04/14/07 13:07 atrig Trigger-Download
  • 199012 04/14/07 13:05 gomez VESDA alert - beamline and air-conditioning
  • 199011 04/14/07 13:00 Elena Kh. L vdc v2?
  • 199009 04/14/07 12:32 Elena Kh. At the end of run 1458 shift tn beam x,y
  • 198998 04/14/07 10:19 Elena Kh. Hall A General Tools and Current Monitor
  • 198997 04/14/07 10:18 Elena Kh. Radiation monitors stripcharts
  • 198996 04/14/07 10:16 Elena Kh. beam info stripchart
  • 198989 04/14/07 08:38 Elena Kh. spot++
  • 198986 04/14/07 08:13 A. Camsonne Owl shift summary
  • 198982 04/14/07 07:35 A. Camsonne Spot ++
  • 198971 04/14/07 04:52 A. Camsonne Radiation monitors stripchart
  • 198970 04/14/07 04:51 A. Camsonne Q1 back
  • 198968 04/14/07 04:47 A. Camsonne Warning message in analyzer
  • 198967 04/14/07 04:44 A. Camsonne Left Q1 trip
  • 198947 04/14/07 00:12 luhj Target Screens
  • 198944 04/14/07 00:04 sawatzky Shift Summary -- SWING April 13, 2007
  • Friday

  • 198942 04/13/07 23:58 sawatzky No ADC data for LHRS S2 and Cerenkov detectors for runs 2431, 2432. Run 2433 looks good
  • 198939 04/13/07 23:19 Makis P spot++ to compare with OTR run
  • 198938 04/13/07 23:16 Mina Kat target Cell_aft alarm
  • 198935 04/13/07 22:46 sawatzky Where to find the target OTR screen
  • 198933 04/13/07 22:29 sawatzky SLI and Target OTR snapshots
  • 198931 04/13/07 22:06 sawatzky Spot++
  • 198928 04/13/07 21:51 sawatzky Incorrect EPICS_CA_ADDR_LIST in '~adaq/.bash_profile'
  • 198923 04/13/07 20:40 sawatzky Target pressure for the last 4 hours
  • 198922 04/13/07 20:38 sawatzky Radiation monitors snapshot
  • 198921 04/13/07 20:37 sawatzky Hall A Tools Snapshot
  • 198920 04/13/07 20:25 Mina Kat Cryotarget controls and He-3 loop screen
  • 198919 04/13/07 20:24 Mina Kat Target Screens
  • 198917 04/13/07 20:12 sawatzky TS11 locked up at end of run 2426
  • 198916 04/13/07 19:28 sawatzky Left Q1 tripped -- reset cleanly
  • 198911 04/13/07 18:58 sawatzky Minor changes to online replay files
  • 198908 04/13/07 18:33 Mina Kat spot++ run 2423
  • 198901 04/13/07 17:04 Mina Kat Target vacuum increase
  • 198898 04/13/07 16:44 sawatzky Alarm Handler wasn't running on hacsbc2 -- started a new process
  • 198894 04/13/07 16:00 R. Michaels day shift summary
  • 198893 04/13/07 15:54 R. Michaels scaler checks
  • 198888 04/13/07 15:00 R. Michaels radmonitors
  • 198885 04/13/07 14:04 LeRose Radiation Monitors
  • 198881 04/13/07 13:29 R. Michaels Q1-L trip
  • 198873 04/13/07 11:46 jpchen target HP heater alarm
  • 198870 04/13/07 11:03 R. Michaels scaler history data in hex
  • 198865 04/13/07 10:08 R. Michaels spot++
  • 198864 04/13/07 10:07 R. Michaels general tools
  • 198863 04/13/07 09:56 L. Zhu more target screens
  • 198862 04/13/07 09:54 L. Zhu target screens
  • 198857 04/13/07 08:27 A. Camsonne Owl shift summary
  • 198856 04/13/07 08:25 A. Camsonne Q1 in good negative polarity starting run 2403
  • 198848 04/13/07 06:54 A. Camsonne Hall A tools
  • 198845 04/13/07 06:41 Chiranjib target screens
  • 198844 04/13/07 06:40 Chiranjib target screens
  • 198829 04/13/07 02:36 Chiranjib Target screens
  • 198828 04/13/07 02:36 Chiranjib target screens
  • 198820 04/13/07 00:38 A. Camsonne Q1 Left trip
  • 198816 04/13/07 00:08 Chiranjib target screens
  • 198815 04/13/07 00:07 Chiranjib target screens
  • 198814 04/13/07 00:01 Solvignon Shift summary
  • Thursday

  • 198810 04/12/07 23:39 A. Camsonne HPH to 0 again
  • 198809 04/12/07 23:15 Makis P Sudden increase in tgt T/heater to 0
  • 198798 04/12/07 20:38 solvignon Spot++: new raster size
  • 198797 04/12/07 20:36 Makis P Raster change
  • 198790 04/12/07 20:01 Solvignon Hall A general tool
  • 198787 04/12/07 19:43 Solvignon Spot++ (zoomed) with the picture this time !
  • 198786 04/12/07 19:41 Solvignon Spot++ (zoomed)
  • 198785 04/12/07 19:28 Solvignon Spot++
  • 198784 04/12/07 19:28 Makis P H3 Loop chart
  • 198781 04/12/07 19:22 Makis P CHL screen
  • 198780 04/12/07 19:22 Makis P target coolant screen
  • 198779 04/12/07 19:21 Makis P target P and T
  • 198776 04/12/07 19:04 Makis P Target Controls screen
  • 198758 04/12/07 16:05 Higinbotham Shift Summary
  • 198757 04/12/07 16:05 L. Zhu target screens
  • 198756 04/12/07 16:04 Higinbotham 100 uA on 20cm 3He
  • 198753 04/12/07 15:05 Higinbotham Beam Back
  • 198750 04/12/07 14:26 Moffit Operational Restrictions..
  • 198749 04/12/07 13:57 spiegel Q1 power supply comm error
  • 198748 04/12/07 13:54 spiegel Q1 power supply comm error
  • 198745 04/12/07 11:20 L. Zhu target to Empty
  • 198744 04/12/07 10:49 Moffit Hall Survey..
  • 198741 04/12/07 08:10 A. Camsonne Owl shift summary
  • 198740 04/12/07 07:56 Elena Kh. radmonitors, HallA General Tools
  • 198739 04/12/07 07:44 A. Camsonne (TECH) Right Q1 error
  • 198738 04/12/07 07:38 A. Camsonne Right Q1 trip
  • 198734 04/12/07 07:14 A. Camsonne Q1 tripped during run 2360
  • 198726 04/12/07 05:58 A. Camsonne Beam position
  • 198722 04/12/07 05:50 A. Camsonne HARP PMT HV tripped
  • 198721 04/12/07 05:45 A. Camsonne Reduced JT Valve for loop1
  • 198720 04/12/07 05:45 Elena Kh. Beamline?
  • 198719 04/12/07 05:39 A. Camsonne Target HPH red alarm
  • 198712 04/12/07 04:41 A. Camsonne Spot 2356
  • 198711 04/12/07 04:34 Elena Kh. HPH ->0 for a few seconds
  • 198708 04/12/07 04:19 Elena Kh. Cryo chart
  • 198707 04/12/07 04:06 Elena Kh. Target screens
  • 198699 04/12/07 02:51 A. Camsonne Spot 2350
  • 198687 04/12/07 00:59 Elena Kh. T7 trigger
  • 198686 04/12/07 00:54 Elena Kh. replay_coin.C
  • 198684 04/12/07 00:40 Elena Kh. replay_coin
  • 198682 04/12/07 00:30 A. Camsonne Beam position oscillating
  • 198679 04/12/07 00:01 E.Chudakov Swing shift summary
  • Wednesday

  • 198676 04/11/07 23:55 Shahinyan Target move
  • 198674 04/11/07 23:20 E.Chudakov spot++ on the dummy
  • 198672 04/11/07 23:08 Shahinyan Target move
  • 198670 04/11/07 22:51 E.Chudakov Tools, beam
  • 198669 04/11/07 22:48 Shahinyan Target
  • 198665 04/11/07 21:57 E.Chudakov Target heater went off
  • 198649 04/11/07 18:30 E.Chudakov Resumed production
  • 198648 04/11/07 18:28 E.Chudakov spot++ on 3He
  • 198646 04/11/07 18:07 Makis P Trigger rates for 85uA empty (no target)
  • 198641 04/11/07 16:16 Shahinyan Target status
  • 198640 04/11/07 16:09 LeRose MicroAmps
  • 198639 04/11/07 16:06 LeRose MicroAmps
  • 198638 04/11/07 16:04 LeRose Shift Summary
  • 198636 04/11/07 15:34 gilman ioc rebooted
  • 198635 04/11/07 15:01 gilman target adjusted from 7.5 to 8.0 K
  • 198634 04/11/07 14:58 LeRose Radiation Monitors
  • 198633 04/11/07 14:27 LeRose Left Q1 is back
  • 198632 04/11/07 14:20 LeRose Left Q1 Trip
  • 198631 04/11/07 13:25 LeRose ep & ARC HV
  • 198630 04/11/07 13:11 LeRose Beam Off
  • 198627 04/11/07 11:47 LeRose Hall A Tools
  • 198612 04/11/07 08:50 Moffit Strange long tail in spot++
  • 198606 04/11/07 08:06 A. Camsonne Owl shift summary 04/11/2007
  • 198605 04/11/07 07:56 A. Camsonne Late entry : additionnal delays on TC stop for
  • 198602 04/11/07 07:49 A. Camsonne Late entry : additionnal delays in left arm
  • 198601 04/11/07 07:42 A. Camsonne Ion chamber strip tool
  • 198600 04/11/07 07:33 A. Camsonne BPM strip chart
  • 198591 04/11/07 05:52 A. Camsonne SPot++ 2315
  • 198588 04/11/07 05:27 Elena Kh. spot++
  • 198585 04/11/07 05:09 A. Camsonne Spot run 2313
  • 198584 04/11/07 05:04 A. Camsonne HARP PMT 2 tripped off
  • 198583 04/11/07 04:57 A. Camsonne Hall A tools
  • 198582 04/11/07 04:56 A. Camsonne spot++ Run2312
  • 198578 04/11/07 04:28 Elena Kh. Target screens
  • 198577 04/11/07 04:25 A. Camsonne Raster size to 5x2 to get 1.5x1.5 on target
  • 198576 04/11/07 04:24 Elena Kh. He3 target
  • 198575 04/11/07 04:18 A. Camsonne Spot++ (-2.25,1) Run 2311
  • 198572 04/11/07 04:11 A. Camsonne Raster with (-2.25,0.8) run 2310
  • 198569 04/11/07 04:04 A. Camsonne Spot++ Run 2309 (-2.25,-0.2)
  • 198566 04/11/07 03:53 A. Camsonne Spot ++ Carbon with hole 2307
  • 198559 04/11/07 03:35 Elena Kh. Carbon Hole target
  • 198558 04/11/07 03:31 A. Camsonne Raster 10x4 mm MCC
  • 198555 04/11/07 03:25 A. Camsonne Raster size seems good 5x2 mm2 MCC gives 1.5x1.5 mm2 at target
  • 198550 04/11/07 03:18 A. Camsonne Spot ++ 2.5x4 MCC
  • 198546 04/11/07 03:07 A. Camsonne Raster with spot++ with 5x4 mm MCC
  • 198543 04/11/07 03:00 A. Camsonne Spot++ at (-2.25,0)
  • 198542 04/11/07 02:59 A. Camsonne Harp scan at good position
  • 198540 04/11/07 02:55 A. Camsonne Beam position : almost perfect !
  • 198534 04/11/07 02:28 Elena Kh. BeO target
  • 198533 04/11/07 02:25 A. Camsonne Harp scan 2341
  • 198530 04/11/07 01:34 A. Camsonne IHA1H03B
  • 198529 04/11/07 01:33 A. Camsonne IHA1H03B Harp odd reading
  • 198528 04/11/07 01:30 A. Camsonne Harp scan 2335-2336
  • 198527 04/11/07 01:20 A. Camsonne MCC can deliver up to 75 uA tonight
  • Tuesday

  • 198513 04/10/07 23:59 sawatzky Shift Summary (Tuesday SWING)
  • 198507 04/10/07 22:43 sawatzky Cryotarget controls snapshots
  • 198506 04/10/07 22:38 sawatzky Target screen snapshots
  • 198505 04/10/07 20:29 sawatzky Hall A Tools snapshot
  • 198504 04/10/07 20:26 sawatzky Target ladder in empty position
  • 198503 04/10/07 19:43 sawatzky ESR 4k flow dip
  • 198502 04/10/07 19:41 sawatzky Left Q1 tripped
  • 198501 04/10/07 19:30 sawatzky Fixes during the controlled access
  • 198500 04/10/07 19:10 sawatzky Target vacuum spikes
  • 198499 04/10/07 19:05 atrig Trigger-Download
  • 198498 04/10/07 19:05 Moffit Another Warning after trigsetup
  • 198497 04/10/07 18:36 sawatzky New encoder positions not quite right yet -- updated positions in text
  • 198496 04/10/07 17:49 sawatzky New step after target IOC reboot: click 'Clear Selected Record' to re-enable motion
  • 198495 04/10/07 17:44 sawatzky Target -> BeO
  • 198494 04/10/07 16:59 sawatzky How to review archived EPICS variables
  • 198493 04/10/07 16:39 sawatzky Going to controlled access
  • 198492 04/10/07 16:04 hyao Target IOC rebooted
  • 198490 04/10/07 15:51 Ole Hansen Day shift summary
  • 198489 04/10/07 14:06 Ole Hansen No beam
  • 198487 04/10/07 11:47 Ole Hansen spot check
  • 198484 04/10/07 11:32 Ole Hansen Hall A tools and current
  • 198481 04/10/07 11:21 A. Camsonne Cerenkov spectrum all T3 run 2272
  • 198478 04/10/07 10:45 Ole Hansen Better beam now
  • 198475 04/10/07 10:42 Ole Hansen Online replay run 2282
  • 198472 04/10/07 09:59 Ole Hansen Beam arrives
  • 198471 04/10/07 09:58 Ole Hansen Harp instructions need update
  • 198469 04/10/07 09:34 hyao Target moved back to He3
  • 198468 04/10/07 09:28 Ole Hansen Beam size ok after all
  • 198467 04/10/07 09:23 hyao Targe Moved to BeO
  • 198466 04/10/07 09:11 Ole Hansen Beam size too big
  • 198465 04/10/07 09:07 Ole Hansen Harps
  • 198464 04/10/07 09:06 Ole Hansen Unable to analyze harp scan
  • 198461 04/10/07 08:05 Makis P, Ele Shift Summary
  • 198458 04/10/07 07:46 Chiranjib Target screens
  • 198457 04/10/07 07:45 Chiranjib target screens
  • 198456 04/10/07 07:33 Elena Kh. Elastic Scattering, 55Q2
  • 198455 04/10/07 07:30 atrig Trigger-Download
  • 198454 04/10/07 07:18 Elena Kh. Cer ped cuts too big?
  • 198453 04/10/07 07:07 Elena Kh. right S1 HV->74% of nominal values
  • 198452 04/10/07 06:15 Elena Kh. tof, run #2272 plots
  • 198451 04/10/07 06:12 Elena Kh. beam will be off for 1, 2 hours
  • 198433 04/10/07 05:08 Elena Kh. Hall A General Tool, Current Monitor
  • 198432 04/10/07 05:06 Elena Kh. run 2272 is not a test
  • 198418 04/10/07 04:42 Chiranjib target moved to He -3
  • 198416 04/10/07 04:26 Elena Kh. spot++
  • 198413 04/10/07 04:17 Elena Kh. spot++
  • 198412 04/10/07 04:15 Makis P harps done, beam of order 100microns
  • 198409 04/10/07 03:39 Makis P. On ID of Run 2263
  • 198408 04/10/07 03:32 Makis P. Harps Scan to begin
  • 198407 04/10/07 03:31 Makis P. Beam is centered
  • 198406 04/10/07 03:25 Makis P. magnet snapshot
  • 198405 04/10/07 03:22 Chiranjib target screens
  • 198404 04/10/07 03:22 Chiranjib target screens
  • 198403 04/10/07 03:16 Makis P. Beam delivery status
  • 198402 04/10/07 02:48 Makis P. Starting getting tuned beam
  • 198400 04/10/07 01:39 Makis P. MCC update: preparing to tune beam for Hall A
  • 198397 04/10/07 00:30 atrig Trigger-Download
  • 198389 04/10/07 00:01 E.Chudakov Swing shift summary
  • Monday

  • 198387 04/09/07 23:56 Shahinyan Target staus
  • 198385 04/09/07 20:34 Shahinyan Target Status
  • 198384 04/09/07 20:08 E.Chudakov Message from PD
  • 198382 04/09/07 18:11 E.Chudakov Target ladder - wrong position
  • 198381 04/09/07 18:07 E.Chudakov Magnets, tools
  • 198380 04/09/07 17:34 E.Chudakov HRSL settings
  • 198379 04/09/07 17:18 LeRose Low Field Magnet Setting
  • 198378 04/09/07 17:17 LeRose Low Field Magnet Setting
  • 198377 04/09/07 17:08 folts tech
  • 198374 04/09/07 16:57 folts right spectrometer
  • 198373 04/09/07 16:55 folts left spectrometer
  • 198372 04/09/07 16:50 Shahinyan Target move
  • 198370 04/09/07 16:09 jpchen Day shift summary
  • 198369 04/09/07 16:06 Adaq Target Status Check
  • 198368 04/09/07 14:23 A. Camsonne Right arm reset fastbus
  • 198366 04/09/07 14:02 Elena Kh. new right S1 HVs
  • 198365 04/09/07 14:00 Elena Kh. Right S1 HVs
  • 198364 04/09/07 13:58 A. Camsonne Detector timing
  • 198363 04/09/07 13:46 jpchen increased T from 7 to 7.5K, reduced flow from 20.5 to 18.5 g/s
  • 198362 04/09/07 13:31 jpchen target charts, T=7K
  • 198361 04/09/07 12:24 A. Camsonne HV left arm
  • 198359 04/09/07 09:07 jpchen tools screen.
  • 198358 04/09/07 09:06 jpchen beam off. energy change
  • 198351 04/09/07 08:15 A. Camsonne Owl shift summary
  • 198347 04/09/07 07:44 A. Camsonne ROC 2 stuck need to be power cycled
  • 198337 04/09/07 05:48 A. Camsonne Spot run 2247
  • 198332 04/09/07 05:15 A. Camsonne Spot run 2245
  • 198329 04/09/07 05:03 A. Camsonne Magnets levels stripchart
  • 198328 04/09/07 04:55 kelleher target: vacuum alarm
  • 198326 04/09/07 04:44 kelleher target motion: 3He target
  • 198324 04/09/07 04:26 A. Camsonne Left Q1 He level
  • 198323 04/09/07 04:15 A. Camsonne (TECH) Fluctuation in He level
  • 198322 04/09/07 04:14 kelleher target alarm and ESR 4k flow.
  • 198315 04/09/07 03:04 A. Camsonne CODA 100% dead time
  • 198310 04/09/07 02:28 A. Camsonne Spot run 2238
  • 198309 04/09/07 02:25 A. Camsonne Beam position strip chart
  • 198308 04/09/07 02:23 A. Camsonne Hall A ion chamber
  • 198306 04/09/07 02:14 kelleher target motion: Dummy Cell
  • 198302 04/09/07 01:37 A. Camsonne General Hall tools
  • 198301 04/09/07 01:37 A. Camsonne Spot ++
  • 198296 04/09/07 01:02 A. Camsonne Scaler map updated
  • 198291 04/09/07 00:18 Mina Kat Shift Summary
  • 198287 04/09/07 00:08 Adaq R Dipole power supply striptool
  • Sunday

  • 198280 04/08/07 23:05 Mina Kat general Accelerator parameters screen
  • 198279 04/08/07 23:01 Mina Kat TS0 crashed, no end run for 2224
  • 198278 04/08/07 22:58 Mina Kat Spot++ run 2228
  • 198277 04/08/07 22:53 Kalyan Target screens
  • 198265 04/08/07 21:17 Mina Kat spot++ run 2222
  • 198256 04/08/07 19:49 Mina Kat spot++ run 2217
  • 198253 04/08/07 19:34 Mina Kat Hall A tools and Current monitor
  • 198248 04/08/07 18:45 Mina Kat spot++ run 2215
  • 198242 04/08/07 18:05 Mina Kat Spot++ for run 2211
  • 198237 04/08/07 17:01 Kalyan Target moved
  • 198236 04/08/07 16:31 folts spectrometer move
  • 198235 04/08/07 16:06 Moffit Hall Survey..
  • 198234 04/08/07 16:06 Moffit Hall Survey..
  • 198232 04/08/07 15:59 Elena Kh. Shift Summary
  • 198231 04/08/07 15:40 Moffit Ion Chamber readings since Friday
  • 198228 04/08/07 15:12 atrig Trigger-Download
  • 198227 04/08/07 15:09 Moffit Momentum set for q2=40
  • 198226 04/08/07 14:55 xqian target movement
  • 198225 04/08/07 14:53 Elena Kh. Additional peak at L vdc tdc, v1
  • 198221 04/08/07 13:58 Elena Kh. At the end of run 2208 TS0 crashed
  • 198220 04/08/07 13:34 R. Michaels re: DT in datamon
  • 198218 04/08/07 13:31 xqian target movement
  • 198213 04/08/07 12:05 Elena Kh. HallA Current Monitor
  • 198206 04/08/07 10:59 Elena Kh. Hall A General Tools
  • 198201 04/08/07 10:22 Elena Kh DT screen
  • 198198 04/08/07 09:52 Elena Kh. DT screen
  • 198193 04/08/07 09:15 Elena Kh. spot++
  • 198192 04/08/07 09:14 Elena Kh. spot++
  • 198184 04/08/07 07:49 suleiman Owl Shift Summary
  • 198163 04/08/07 03:58 A. Camsonne Run 2184 Coda crashed ROC1
  • 198161 04/08/07 03:50 A. Camsonne Run 2178 2179, R coinc time +24 ns
  • 198152 04/08/07 02:31 suleiman spot++
  • 198150 04/08/07 02:17 atrig Trigger-Download
  • 198148 04/08/07 02:14 A. Camsonne R arm VDC efficiency
  • 198145 04/08/07 01:47 suleiman Accelerator General Screen
  • 198143 04/08/07 01:42 mina/suleima Moved to Q2=34 fm-2
  • 198142 04/08/07 01:17 A. Camsonne Target screens
  • 198141 04/08/07 01:00 atrig Trigger-Download
  • 198139 04/08/07 00:40 suleiman Helium Q2=25 fm-2, H-arm at -2.0%
  • 198136 04/08/07 00:12 suleiman Helium Q2=25 fm-2, H-arm at +1.02%
  • 198134 04/08/07 00:03 Mina Kat Shift Summary
  • Saturday

  • 198131 04/07/07 23:42 Mina Kat spot++ dummy target run 2175
  • 198130 04/07/07 23:39 Makis P dead time discrepancy?
  • 198129 04/07/07 23:36 Mina Kat spot++ for run 2173
  • 198128 04/07/07 23:35 xqian Change target to Al
  • 198127 04/07/07 23:34 Mina Kat spot++ for run 2174
  • 198122 04/07/07 22:52 Moffit Target and Magnets stable, keeping CHL 20.5 g/s to target for now.
  • 198119 04/07/07 22:31 folts Q1/tech
  • 198110 04/07/07 21:11 Mina Kat General Tools and beam current monitor
  • 198108 04/07/07 21:05 Mina Kat spot++
  • 198107 04/07/07 21:04 Xqian Target is stable
  • 198104 04/07/07 20:52 atrig Trigger-Download
  • 198103 04/07/07 20:07 R. Michaels re: trigger download error
  • 198102 04/07/07 19:07 Mina Kat Right Q1 trips again
  • 198101 04/07/07 17:35 Mina Kat Hall Access Right Q1 trip
  • 198100 04/07/07 17:24 Moffit Right arm S1 adc spectra... HV too high
  • 198097 04/07/07 16:52 xqian Open JT valve
  • 198096 04/07/07 16:50 Moffit Suspicious S1 R-PMT ADC spectrum
  • 198091 04/07/07 16:16 bogdanw shift summary
  • 198088 04/07/07 16:13 atrig Trigger-Download
  • 198086 04/07/07 16:12 Elena Kh. Trigger Download ?
  • 198085 04/07/07 16:10 bogdanw Tools for He elastic
  • 198082 04/07/07 15:46 Elena Kh. Loop1 Fill_Target_R
  • 198081 04/07/07 15:39 atrig Trigger-Download
  • 198080 04/07/07 15:38 atrig Trigger-Download
  • 198079 04/07/07 15:35 atrig Trigger-Download
  • 198078 04/07/07 15:33 atrig Trigger-Download
  • 198077 04/07/07 15:29 Adaq HRS-R HV for He run
  • 198074 04/07/07 14:27 R. Michaels re: low ADC on L-arm
  • 198070 04/07/07 13:31 Elena Kh. Target screens
  • 198051 04/07/07 10:13 Elena Kh. Removed adc cuts for the shower detector
  • 198050 04/07/07 10:03 bogdanw Left arm HV
  • 198041 04/07/07 08:45 Adaq beam line HV
  • 198030 04/07/07 07:50 suleiman Owl Shift Summary
  • 198029 04/07/07 07:40 suleiman Rates at 20 uA on He3
  • 198024 04/07/07 07:18 Adaq General Accelerator Parameters Screen
  • 198022 04/07/07 07:16 A. Camsonne Coincidence time good.
  • 198020 04/07/07 07:09 suleiman Hall A General Tools
  • 198019 04/07/07 07:08 atrig Trigger-Download
  • 198014 04/07/07 06:13 A. Camsonne Target screens
  • 198013 04/07/07 06:12 A. Camsonne Target screens
  • 198011 04/07/07 05:46 suleiman Rates at Raster 10.0 by 10.0
  • 198008 04/07/07 05:17 suleiman Helium target moved up 2 mm
  • 198007 04/07/07 05:11 atrig Trigger-Download
  • 198005 04/07/07 04:51 suleiman Helium target moved down 2 mm
  • 198002 04/07/07 04:41 suleiman Rates at Raster 3.5 by 2.0
  • 198001 04/07/07 04:37 suleiman Rates at Raster 10.0 by 10.0
  • 198000 04/07/07 04:31 suleiman Rates at Raster 10.0 by 2.0
  • 197993 04/07/07 03:46 suleiman L-arm Cerenkov Signals are low
  • 197987 04/07/07 02:47 suleiman Keep the old HV till the morning
  • 197985 04/07/07 02:36 suleiman L-arm S1 HV before the change
  • 197982 04/07/07 02:33 suleiman L-arm S1/S2 ADCs
  • 197978 04/07/07 02:13 atrig Trigger-Download
  • 197974 04/07/07 01:55 A. Camsonne Left S1R TDC missing
  • 197971 04/07/07 01:43 A. Camsonne Added 56 ns to TC right arm
  • 197967 04/07/07 01:29 atrig Trigger-Download
  • 197955 04/07/07 00:32 A. Camsonne Moving to He3 target Loop1
  • 197954 04/07/07 00:32 suleiman Raster Screen - Final
  • 197953 04/07/07 00:29 Adaq Beam Position - Final
  • 197952 04/07/07 00:27 Adaq Beam Spot++ - Final
  • 197947 04/07/07 00:23 A. Camsonne Target moved to BeO
  • Friday

  • 197933 04/06/07 23:56 Chiranjib IOC reboot
  • 197932 04/06/07 23:50 Chiranjib target moved to Carbon with hole
  • 197931 04/06/07 23:50 Ole Hansen Shift summary
  • 197929 04/06/07 23:29 Ole Hansen Target offsets
  • 197928 04/06/07 23:03 Ole Hansen Spot size
  • 197926 04/06/07 23:00 Ole Hansen spot++ on BeO
  • 197925 04/06/07 22:57 Chiranjib target moved to Be0
  • 197923 04/06/07 22:46 Ole Hansen Optimized beam position on carbon
  • 197919 04/06/07 22:25 Ole Hansen Runs 2103/2104 junk
  • 197910 04/06/07 22:03 Ole Hansen Raster pattern run 2101 - hole centered
  • 197908 04/06/07 21:52 Ole Hansen Raster pattern run 2100
  • 197906 04/06/07 21:44 Ole Hansen Carbon hole found
  • 197905 04/06/07 21:42 Ole Hansen spot check run 2100 bpm x/y=-1/+3
  • 197900 04/06/07 21:01 Ole Hansen spot++ 3x3 bpm x/y=0/1
  • 197897 04/06/07 20:58 Ole Hansen spot++ for x/y=0/3
  • 197896 04/06/07 20:55 Ole Hansen Spot check ratser 6x6
  • 197891 04/06/07 20:40 Ole Hansen Spot check
  • 197885 04/06/07 18:55 atrig Trigger-Download
  • 197881 04/06/07 18:35 Chiranjib target moved to Be0
  • 197866 04/06/07 17:39 A. Camsonne Spot Run 2086
  • 197865 04/06/07 17:31 atrig Trigger-Download
  • 197864 04/06/07 17:28 Chiranjib target moved to carbon hole
  • 197863 04/06/07 17:24 A. Camsonne Harp scan 2332
  • 197857 04/06/07 16:21 Makis P. Shift Summary
  • 197850 04/06/07 15:22 A. Camsonne Harp scan 2331
  • 197847 04/06/07 14:54 jpchen target cooled-down and condition
  • 197839 04/06/07 13:47 A. Camsonne S1 right pedestal check
  • 197838 04/06/07 13:44 A. Camsonne Pedestal subtraction on
  • 197836 04/06/07 13:43 A. Camsonne S2m pedestal check
  • 197835 04/06/07 13:19 hyao Target = Empty
  • 197834 04/06/07 13:19 A. Camsonne No pedestal subtraction for next runs starting 2076
  • 197831 04/06/07 12:22 Moffit Hall Lockup
  • 197829 04/06/07 12:10 spiegel Pre-Beam Checklist
  • 197825 04/06/07 11:27 Jack Detector configuration
  • Thursday

  • 197807 04/05/07 20:23 gomez E04-018 Left/Right HRS momentum set for (e,e'p) checkout; P_L = 0.457 GeV/c, P_R = 0.899 GeV/c
  • 197796 04/05/07 16:57 Elena Kh. changed bcm1,2 coefficient
  • 197795 04/05/07 16:54 Elena Kh. mcc: bcm1,2 plots
  • 197794 04/05/07 16:45 Elena Kh. Hall A Current Monitor
  • Wednesday

    
    

    Tuesday

  • 197769 04/03/07 22:45 A. Camsonne Sweep aborted
  • 197768 04/03/07 21:33 Adaq Hall A tools
  • 197759 04/03/07 17:03 Jack Ethanol level in bubbler for VDC gas mixing system
  • 197758 04/03/07 16:31 M. Stevens Rt. Dipole new failure mode
  • 197757 04/03/07 15:06 Moffit Hall Lockup after 6pm today
  • Monday

  • 197756 04/02/07 11:27 Moffit Shifts cancelled through Wednesday Day
  • 197755 04/02/07 11:13 atrig Trigger-Download
  • 197749 04/02/07 09:39 atrig Trigger-Download
  • 197748 04/02/07 09:39 A. Camsonne DAQ in running mode
  • 197747 04/02/07 09:22 fansler Left and Right Q1's
  • Sunday

  • 197746 04/01/07 23:21 atrig Trigger-Download
  • 197740 04/01/07 14:58 R.M./A.C. trigger mods part III
  • 197739 04/01/07 14:39 atrig Trigger-Download
  • 197738 04/01/07 14:35 atrig Trigger-Download
  • 197737 04/01/07 14:30 atrig Trigger-Download
  • 197736 04/01/07 14:20 atrig Trigger-Download
  • 197735 04/01/07 14:16 atrig Trigger-Download
  • 197734 04/01/07 14:15 atrig Trigger-Download
  • 197733 04/01/07 14:15 atrig Trigger-Download
  • 197732 04/01/07 14:14 atrig Trigger-Download
  • 197731 04/01/07 14:12 atrig Trigger-Download
  • 197730 04/01/07 14:11 atrig Trigger-Download
  • 197729 04/01/07 14:11 atrig Trigger-Download
  • 197728 04/01/07 14:10 atrig Trigger-Download
  • 197727 04/01/07 14:05 atrig Trigger-Download
  • 197726 04/01/07 14:04 atrig Trigger-Download
  • 197725 04/01/07 14:04 atrig Trigger-Download
  • 197724 04/01/07 14:01 atrig Trigger-Download
  • 197723 04/01/07 14:00 atrig Trigger-Download
  • 197722 04/01/07 13:58 atrig Trigger-Download
  • 197721 04/01/07 13:58 atrig Trigger-Download
  • 197720 04/01/07 13:57 atrig Trigger-Download
  • 197719 04/01/07 13:51 atrig Trigger-Download
  • 197718 04/01/07 13:51 atrig Trigger-Download
  • 197717 04/01/07 13:50 atrig Trigger-Download
  • 197716 04/01/07 13:47 atrig Trigger-Download
  • 197715 04/01/07 13:44 atrig Trigger-Download
  • 197714 04/01/07 13:41 atrig Trigger-Download
  • 197713 04/01/07 13:40 atrig Trigger-Download
  • 197712 04/01/07 13:36 R. Michaels/ RT & HV fixes
  • 197704 04/01/07 13:11 atrig Trigger-Download
  • 197702 04/01/07 12:52 atrig Trigger-Download
  • 197698 04/01/07 12:06 R. Michaels L-arm RT problems, HV problem
  • 197692 04/01/07 10:45 atrig Trigger-Download
  • 197691 04/01/07 10:04 atrig Trigger-Download